OpenStudy (samsan9):

if x-1 is a factor of P(x)=x^3-5x^2+7x-3, which of the following represents the complete factorization for P(x)?

OpenStudy (samsan9):

A.) (x-1)(x+3)(x+1) B.) (x-3)(x-1)(x-1) C.) (x-3)(x+3)(x-1) D.) (x-3)((x-1) (x+1)

OpenStudy (anonymous):

you may divide x^3-5x^2+7x-3 by x-1 and compare the quotient

OpenStudy (anonymous):

other wise use factor remainder theorem

OpenStudy (anonymous):

or you may factorise completetly

OpenStudy (samsan9):

which one is easier?

OpenStudy (anonymous):

using division

OpenStudy (anonymous):

but i may show u by factoring

OpenStudy (anonymous):

x^3-5x^2+7x-3 =x^3 -x^2 -4x^2 +4x +3x - 3 =x^2(x-1)-4x(x-1) +3(x-1) =(x-1)(x^2-4x+3)

OpenStudy (anonymous):

=(x-1)(x^2-4x+3) =(x-1)(x^2 -x-3x+3) =(x-1)(x(x-1)-3(x-1)) =(x-1)(x-1)(x-3)

OpenStudy (samsan9):

so B?

OpenStudy (anonymous):

at least u should be sure of that ...

OpenStudy (samsan9):

thanks :D

OpenStudy (anonymous):

yw